Circular No. 8832

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V5558 SAGITTARII = NOVA SAGITTARII 2007 S. Nakano, Sumoto, Japan, reports the discovery by Yukio Sakurai

(Mito, Ibaraki-ken) of a possible nova (mag 10.3) on two 20-s CCD exposures taken on Apr. 14.777 UT with a Fuji FinePix S2 Pro Digital camera (+ Nikon 180-mm f/2.8 lens), giving the position for the new object as α = 18^{h}10^{m}18^{s}.4, δ = −18^{o}46′51′′ (equinox 2000.0). Sakurai adds that nothing is visible at this location on an image taken on Apr. 9.8 (limiting magnitude 11.4). Nakano provides the following position end figures for his own measurement of Sakurai’s discovery ‘jpeg’ image: 18^{s}.47, 52′′.0. On Apr. 19.712, Nakano obtained his own unfiltered CCD frames of the new object (at mag 10.2) with a 0.25-m f/4.8 reflector, providing the position end figures 18^{s}.19, 50′′.9 (but apparently involved with a nearby 12th-magnitude star). Nakano adds that K. Itagaki (Yamagata, Japan) obtained an unfiltered CCD image on Apr. 19.745 showing the new object at mag 9.8, with Itagaki measuring position end figures 18^{s}.27, 52′′.1; Itagaki adds that the closest star on the same frame has position end figures 18^{s}.03, 46′′.1 (mag 12.1). Neither the USNO catalogue nor the Digitized Sky Survey indicates any precursor at the position of the new star. R. Koff (Bennett, CO, U.S.A.) reports the following position end figures 18^{s}.21, 51′′.8 and magnitude V = 9.8 for the apparent nova from CCD images obtained with a 0.25-m f/10 reflector on Apr. 20.4, adding that it is blended with the star GSC 6272-0182, which is just northwest of the new object.

H. Yamaoka, Kyushu University, writes that, following posting of this object on the Central Bureau’s unconfirmed-objects webpage, K. Haseda (Toyohashi, Aichi, Japan) has located several prediscovery images of the new star with a Canon EOS-5D digital camera (+ 120-mm-f.l. f/3.5 lens), providing the following estimated magnitudes: Apr. 7.780 UT, [11.8; 11.792, 11.2 (correction to date given on CBET 931); 13.793, 10.8; 14.777, 10.4. Yamaoka adds that Y. Nakamura (Kameyama, Mie, Japan) obtained an unfiltered CCD image with a 135-mm-f.l. f/2.8 camera lens on Apr. 19.803 that showed the apparent nova at mag 9.9, and the new star was evidently not present on an image taken by Nakamura on Apr. 9.792.

N. N. Samus, Institute of Astronomy, Russian Academy of Sciences, informs us that this nova has been given the designation V5558 Sgr.

2007 April 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green